OpenStudy (anonymous):

solve (1/36)=6^(x-3)

OpenStudy (anonymous):

take the log of both sides log(1/36)=log(6^(x-3))

OpenStudy (anonymous):

so then its -1.556=46.689

OpenStudy (anonymous):

the answer is 1

OpenStudy (anonymous):

oh thanks :D

OpenStudy (anonymous):

log1-log(36)=(x-3)log(6) log1=0 -log(36)/log(6)=x-3 -2=x-3 x=1